OpenStudy (anonymous):

Graph the linear inequality: x-y>3

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=x-y%3E3

OpenStudy (anonymous):

Did you get the answer?

OpenStudy (anonymous):

Yes, I did. But I can't understand the way it was graphed.

OpenStudy (anonymous):

Okay what is the equation your using to graph? Rem you have to set it to y> in order to graph

OpenStudy (anonymous):

=x-y>3 =-y>3+x =y>(3+x)/(-1) Plug in positive and neg values for x. So by hand construct a table using -5, -2, 0, 2, 5 any x-values that are whole numbers

OpenStudy (anonymous):

Or use a graphing calc